OpenStudy (anonymous):

Question: If the latest planet is 600 Light years away and a light year is roughly 6 trillion miles that would be 3600 Trillion miles away? How fast would you need to travel to make it there in 5 80 year life spans?

OpenStudy (anonymous):

3,600,000,000,000,000 / 3,504,000 hrs = mph - That look right?

OpenStudy (anonymous):

1,027,397,260.3 - rounded?

OpenStudy (anonymous):

You can not travel 600 light years in 400 years.

OpenStudy (anonymous):

You need to go 1.5 time the speed of light... that's not possible

OpenStudy (turingtest):

\[600(6\times10^{12})=6\times10^{14}=distance\]in miles\[5\times80\times365\times24=time\]in hrs divide that out, whatever it is...

OpenStudy (turingtest):

I got \[1.71\times10^8mph\]

OpenStudy (turingtest):

\[6.71\times10^8\]miles per hour is the speed of light, so I don't see how henkjan says that you need to travel faster that the speed of light...

OpenStudy (anonymous):

Something is 600 light years away (the distance a lightbeam would travel in 600 years). If you gonna make this in 400 years, you need to go FASTER than light.

OpenStudy (anonymous):

I got 1005924943.5 mph

OpenStudy (turingtest):

right... so where did I go wrong?

OpenStudy (anonymous):

In reply to henkan - I just wanted a reference to the Notable mile per hour to create a reference to the speed it would take. Using the light speed limitation I would be curious as to the time it would take at max Light speed

OpenStudy (anonymous):

600 lightyears with the speed of light? That gonna cost you 600 years mate (H)

OpenStudy (turingtest):

yes, that seems to be the obvious way to look at it, how did I not see that?

OpenStudy (turingtest):

oh duh \[600\times(6\times10^{12})=36\times10^{14}\]which leads to a speed of about \[10^{9}mi/hr\]which is about 1.5 times the speed of light. I just forgot to multiply the 6's :P but henkjan's way is the smart way, so go with that.

OpenStudy (jamesj):

Now a fun variant for those who know special relativity. How fast would you have to travel to arrive there in 10 years, from the perspective of the space travelers? How long would that trip appear to take on earth?

OpenStudy (anonymous):

slowdown of time: 1/(sqrt(1-(v/c)^2) traveltime without slowdown effect: 600/(v/c) combined: 600/(c/v)*(sqrt(1-(v/c)^2)= 10 v = 0.99986*c on earth it would look like 600/0.99986 = 600.08 years